Two sample drawn from two different populations

Mean Standard deviation Size

Sample- I 165 16 65

Sample- II 150 15 52


Test at 5% level of significance that the difference of mean is 12


1
Expert's answer
2021-05-07T09:27:06-0400

Two sample t-test assuming equal variances.

Null hypothesis "H_0:\mu_1=\mu_2" .

Alternative hypothesis "H_a:\mu\ne\mu_2 ."

Pooled variance: "s^2=\frac{64*16^2+51*15^2}{65+52-2}=171.82."

Test statistic: "\frac{165-150-12}{\sqrt{171.82(\frac{1}{65}+\frac{1}{53})}}=1.04."

Degrees of freedom: "df=65+52-2=115."

P-value: "p<0.3024."

Since the P-value is greater than 0.05, fail to reject the null hypothesis.

There is no sufficient evidence that the population difference of means is not 12.